# Anisotropic flow fluctuation as a possible signature of clustered nuclear geometry in O–O collisions at the Large Hadron Collider

### v1

@@ Abstract
Nuclei having $4n$ number of nucleons are theorized to possess clusters of $\alpha$ particles (4He nucleus). The Oxygen nucleus (16O) is a doubly magic nucleus, where the presence of an $\alpha$-clustered nuclear structure grants additional nuclear stability. In this study, we exploit the anisotropic flow coefficients to discern the effects of an $\alpha$-clustered nuclear geometry w.r.t. a Woods-Saxon nuclear distribution in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using a hybrid of IP-Glasma + MUSIC + iSS + UrQMD models. In addition, we use the multi-particle cumulants method to measure anisotropic flow coefficients, such as elliptic flow ($v_{2}$) and triangular flow ($v_{3}$), as a function of collision centrality. Anisotropic flow fluctuations, which are expected to be larger in small collision systems, are also studied for the first time in O–O collisions. It is found that an $\alpha$-clustered nuclear distribution gives rise to an enhanced value of $v_{2}$ and $v_{3}$ towards the highest multiplicity classes. Consequently, a rise in $v_{3}/v_{2}$ is also observed for the (0-10)centrality class. Further, for $\alpha$-clustered O–O collisions, fluctuations of $v_{2}$ are larger for the most central collisions, which decrease towards the mid-central collisions. In contrast, for a Woods-Saxon 16O nucleus, $v_{2}$ fluctuations show an opposite behavior with centrality. This study, when confronted to experimental data may reveal the importance of nuclear density profile on the discussed observables.
@@ I Introduction
The primary goal of relativistic heavy-ion collisions at the RHIC and the LHC is to create a deconfined thermalised medium of partons, also known as Quark Gluon Plasma (QGP), which is, by our understanding, have existed a few microseconds after the Big Bang. However, due to the very low lifespan of QGP, it can not be observed directly in experiments. Instead, several indirect signatures from experimentally measurable observables can signify the presence of such a deconfined strongly interacting matter in heavy-ion collisions. One such observable is azimuthal anisotropy, which characterises the collective behaviour of the medium formed in heavy-ion collisions. It is quantified in terms of the Fourier expansion coefficients of azimuthal distribution of final state charged hadrons, as follows [1]:
@@
Here, $\phi$ is the azimuthal angle, $\psi_{n}$ stands for the $n$th order symmetry plane angle according to the give $v_{n}$ coefficients of the Fourier expansion. Here, $v_{2}$ quantifies the elliptic flow, $v_{3}$ gives the estimates for triangular flow, and so on. Both $v_{2}$ and $v_{3}$ depend upon the equation of state and transport properties of the QCD medium formed during heavy-ion collisions. In addition, studies based on hydrodynamical calculations show that the flow coefficients are affected strongly by a change in the ratio of shear viscosity to entropy density ($\eta/s$[2, 3]. Moreover, the sensitivity of flow coefficients to $\eta/s$ is large for higher-order harmonics, i.e., $v_{3}$ is more sensitive to $\eta/s$ as compared to $v_{2}$ [3]. In several MC studies, $v_{2}$ and $v_{3}$ are found to be associated with the initial eccentricities of the participating nucleons, where the values of anisotropic flow coefficients can change by modifying the nuclear distribution [6, 4, 5]. In addition, in Xe–Xe collisions, an observation of a large value of $v_{2}$ in most central collisions, as compared to that observed in Pb–Pb collisions, is anticipated due to a deformed nuclear structure of Xe nucleus [7, 8, 10, 9]. In a similar line, the nuclear deformation of Uranium, being a prolate shape nucleus is reflected in a higher $v_{2}$, when compared to that is observed in the most central collisions of Au nuclei, which is close to a spherical nucleus [11]. Similarly, one notices that the values of $v_{2}$ and $v_{3}$ vary with the change in the deformation parameters of Uranium in U–U collisions, and the effect is enhanced for $v_{2}$, as compared to $v_{3}$ [4, 5, 12].
@@
Figure 1: Pictorial representation of an $\alpha$-clustered (left) and non-clustered (right) distribution of nucleons inside the oxygen nucleus.
@@
Elliptic flow is one of the observables that can probe the initial state of heavy-ion collisions [13]. However, the value of elliptic flow fluctuates from one event to another, which reflects the fluctuation in the initial energy density and event-by-event participant nucleon distribution before QGP is formed [14]. These fluctuations in the initial density profile of the participating nucleons can lead to fluctuations in the even harmonic flow coefficients and generate odd flow harmonics [14]. Further, the fluctuations in the initial energy density can develop symmetry planes of different orders in various kinematic regions [15, 16, 17]. In other words, the initial density fluctuations can develop transverse momentum ($p_{\rm T}$) and pseudorapidity ($\eta$) dependence of $\psi_{n}$. Therefore, to comprehend the event-by-event initial density fluctuations and their impact on medium evolution, the study of anisotropic flow fluctuations becomes crucial. In both experimental and theoretical frontiers, a few studies on anisotropic flow fluctuations have been performed in heavy-ion collisions [18, 19, 20, 21]. The Pb–Pb collisions at the LHC energies hint at non-Gaussian initial density fluctuations, and consequently, it led the heavy-ion physicists to constrain their probability distribution functions [22, 23]. However, the studies of anisotropic flow fluctuations in small collision systems are not yet well understood. Moreover, one may conceive that for a small number of participants, the contribution to fluctuations can increase. Finally, if the evolution of anisotropic flow fluctuations has more than one contributor, then the results can vary significantly from the expectations.
@@
In addition to the collective behaviour of the QGP, several other signatures for the presence of a QGP medium have been observed in heavy-ion collisions. For the measurement of such signatures, the baseline measurements are obtained from proton-proton ($pp$) collisions, where the formation of a QCD medium is not anticipated. However, recent observations of strangeness enhancement [24], ridge-like structure [26, 25], radial flow effects [27, 28, 29], etc., in high multiplicity $pp$ collisions have highlighted the significance of studying the small collision systems from different aspects. Recently, LHC and RHIC have a plan to inject Oxygen nuclei (16O) to perform p-O and O–O collisions [30, 31], which can bridge the multiplicity gap between $pp$, p-Pb and Pb–Pb collisions and may provide insight into the possible formation of QGP droplets in small systems. Furthermore, Oxygen is a stable nucleus having a double magic number and is theorized to possess clusters of $\alpha$ (4He) particles, where the $\alpha$-particles arrange themselves in the corners of a tetrahedron [32, 33, 34, 35, 36, 37, 38]. Figure 1 shows a pictorial representation of the nucleons inside an oxygen (16O) nucleus for an $\alpha$-clustered structure (left) and non-clustered distribution of nucleons (right). In recent years, there have been several studies performed to understand the O–O collisions both at RHIC and LHC energies [44, 45, 43, 39, 40, 41, 6, 42, 46]. A few of them investigate the use of hydrodynamical models [47, 48, 49], and Glauber model calculations [39, 40, 41]. Some intend to study the parton energy loss and jet quenching effects in O–O collisions [50, 51]. Interestingly, there have been a few phenomenological studies that aim to establish the effect of the presence of an $\alpha$-cluster nucleus in the final state particle production [44, 45, 43, 53, 54, 55, 56, 6, 52, 57, 58, 59]. A few of these studies employ O–O collisions, while others employ O-Au, C-Au, O-Pb, and Ne-Pb collisions. Nevertheless, the presence of a clustered nuclear structure of 16O is reflected in final state azimuthal anisotropy when compared to a Woods-Saxon nuclear density profile. In Ref. [52], using a multi-phase transport model, authors have shown that an away side broadening is observed in the two-particle correlation function for an $\alpha$-clustered particle. In addition, an increased value of $v_{3}$, and $v_{3}/v_{2}$ for most central O–O collisions are also reported for an $\alpha$-clustered structure. A similar observation was reported in Ref. [44], where authors study the anisotropic flow in C+Au collisions with a clustered structure of 12C nucleus. However, so far, no studies have shown the effect of change in nuclear density profile on event-by-event flow fluctuations. In this paper, for the first time, we report the behaviour of azimuthal anisotropy and flow fluctuations in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using a hydrodynamically expanding system. For this study, we use a hybrid of IPGlasma initial condition, MUSIC hydrodynamics, iSS as a particlisation and UrQMD as an afterburner for the hadronic cascade. Here, we compared results from an $\alpha$-clustered nuclear distribution with a Woods-Saxon nuclear distribution.
@@ II.1 Simulation framework
In this work, the simulation framework implements the impact-parameter-dependent Glasma (IP-Glasma) initial conditions, MUSIC relativistic viscous hydrodynamics, particle sampling using the iSS package, and performs hadronic cascade through the UrQMD transport model. This model provides a good description of the particle production and flow in heavy-ion collisions [60]. The key aspects of these models are also discussed briefly below, including the parameters and settings necessary to reproduce the results reported in this study.
@@ II.1.1 IP-Glasma: Initial condition
The IP-Glasma model is used to describe the dynamics of the gluon fields during the collisions of nuclear matter at relativistic speeds [61, 62]. It provides the initial energy-momentum tensor of the classical Yang-Mills (CYM) color Glasma fields ($T_{\rm YM}^{\mu\nu}$) which can be further evolved using relativistic hydrodynamics. This model is based on the color glass condensate effective theory [63, 64] and uses a classical description of gluon production. This model includes fluctuations in the distributions of nucleons in the nuclear wave functions and also fluctuations in the color charge distributions inside the nucleons. The color charge fluctuations are modeled via the impact-parameter-dependent dipole saturation model (IP-Sat) [65, 66]. For the sub-nucleonic fluctuations, three hot spots are introduced in the proton thickness function, and each hot spot is parametrized using a Gaussian distribution in the transverse plane. The spatial positions of the nucleons inside the colliding nuclei are obtained from the Woods-Saxon distribution function as given below,
@@
Here, $\rho(r)$ is the nuclear charge density at a radial distance $r$ from the center of the nucleus. $r_{0}$, $w$, and $a$ are the mean radius of the nucleus, the nuclear deformation parameter, and nuclear skin depth, respectively. For 16O nucleus, $r_{0}=2.608$ fm, $a=0.513$ fm, $w=-0.051$ [67]. In addition, the exotic $\alpha$–cluster geometry for the 16O nucleus has also been implemented in this study. Four such $\alpha$–particles ($2p,~{}2n$) are placed at the vertices of a regular tetrahedron with side length 3.42 fm, which makes the root mean square (RMS) radius of 16O nucleus to 2.699 fm [6, 42, 43]. The spatial positions of the nucleons inside an $\alpha$–particle are sampled using the Woods-Saxon distribution given in Eq. (2), with the parameters chosen for the 4He nucleus, which are $r_{0}=0.964$ fm, $w=0.517$ and $a=0.322$ fm. This gives the RMS radius of 1.676 fm for the $\alpha$–particle. A minimum separation distance of 0.4 fm between the nucleons is imposed. Finally, to randomize each nucleus, the entire tetrahedral configuration is rotated through its center randomly in the polar and azimuthal coordinates before its initialization in IP-Glasma. Figure 2 shows the nucleon distribution of 16O nuclei in the x-y plane following $\alpha$-cluster (left) and Woods-Saxon (right) nuclear density profiles. In Fig. 2, for the $\alpha$-cluster nuclear density profile of 16O nuclei, the tetrahedral configuration is not rotated for better visualization of the clustered geometry.
Figure 2: The nucleon distribution of 16O nuclei in the x-y plane following $\alpha$-cluster (left) and Woods-Saxon (right) nuclear density profiles plotted with one million nuclei in each case.
@@ II.1.2 MUSIC: Hydrodynamics
Under the assumption of system under local thermal equilibrium, the initial energy-momentum tensor obtained from the IP-Glasma model at thermalization time $\tau_{0}=0.4$ fm is evolved using the MUSIC framework of relativistic hydrodynamics following the conservation law $\partial_{\mu}T^{\mu\nu}=0$ [70, 71, 72]. The energy-momentum tensor is composed of both ideal and dissipative parts, which are combined to give, $T^{\mu\nu}=eu^{\mu}u^{\nu}-(P+\Pi)\Delta^{\mu\nu}+\pi^{\mu\nu}$. Here, $e$ is the energy density, $u^{\mu}$ is the fluid flow four-velocity, $P$ is the thermodynamic pressure, $\Pi$ is the bulk viscous pressure, and $\pi^{\mu\nu}$ is the shear viscous tensor. Again, the projection operator is defined as $\Delta^{\mu\nu}=g^{\mu\nu}-u^{\mu}u^{\nu}$. In this study, we use the boost invariant picture of hydrodynamics and use the equation-of-state parametrization "s95p-v1.2", which is obtained from the interface of lattice data and hadron resonance gas [73]. We use a fixed $\eta/s=0.12$, and a temperature dependent $\zeta/s(T)$ as described in Ref. [49]. In MUSIC, the hydrodynamic equations are solved using the Kurganov-Tadmor (KT) algorithm [74, 75].

### v2

@@ Abstract
Nuclei having $4n$ number of nucleons are theorized to possess clusters of $\alpha$ particles (4He nucleus). The Oxygen nucleus (16O) is a doubly magic nucleus, where the presence of an $\alpha$-clustered nuclear structure grants additional nuclear stability. In this study, we exploit the anisotropic flow coefficients to discern the effects of an $\alpha$-clustered nuclear geometry with respect to a Woods-Saxon nuclear distribution in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using a hybrid of IP-Glasma + MUSIC + iSS + UrQMD models. In addition, we use the multi-particle cumulants method to measure anisotropic flow coefficients, such as elliptic flow ($v_{2}$) and triangular flow ($v_{3}$), as a function of multiplicity class. Anisotropic flow fluctuations, which are expected to be larger in small collision systems, are also studied for the first time in O–O collisions. It is found that an $\alpha$-clustered nuclear distribution gives rise to an enhanced value of $v_{2}$ and $v_{3}$ for the low-multiplicity events. Consequently, a rise in $v_{3}/v_{2}$ is also observed for the 0–10% multiplicity class. Further, for $\alpha$-clustered O–O collisions, fluctuations of $v_{2}$ are larger for the highest multiplicity events, which decrease as the final-state multiplicity decreases. In contrast, for a Woods-Saxon 16O nucleus, $v_{2}$ fluctuations show an opposite behavior with decreasing multiplicity. When confronted with experimental data, this study may reveal the importance of the nuclear density profile on the discussed observables and provide physics validation for the hybrid model discussed in this work.
@@ I Introduction
The primary goal of relativistic heavy-ion collisions at the RHIC and the LHC is to create a deconfined thermalized medium of partons, also known as Quark Gluon Plasma (QGP), which is believed to have existed a few microseconds after the Big Bang. However, due to the transient nature of QGP, it can not be observed directly in experiments. Instead, several indirect signatures from experimentally measurable observables can signify the presence of such a deconfined strongly interacting matter in heavy-ion collisions. One such observable is azimuthal anisotropy, which characterizes the collective behavior of the medium formed in heavy-ion collisions. It is quantified in terms of the Fourier expansion coefficients of azimuthal distribution of final state charged hadrons, as follows [1]:
@@
Here, $\phi$ is the azimuthal angle, $\psi_{n}$ stands for the $n$th order symmetry plane angle and $v_{n}$ being the $n$th order flow coefficients. Here, $v_{2}$ quantifies the elliptic flow, $v_{3}$ gives the estimate for triangular flow, and so on. Both $v_{2}$ and $v_{3}$ depend upon the transport properties of the QCD medium formed during heavy-ion collisions. In addition, studies based on hydrodynamic calculations show that the flow coefficients are affected strongly by a change in the ratio of shear viscosity to entropy density ($\eta/s$[2, 3]. Moreover, the sensitivity of flow coefficients to $\eta/s$ is large for higher-order harmonics, i.e., $v_{3}$ is more sensitive to $\eta/s$ as compared to $v_{2}$ [3]. In several studies, $v_{2}$ and $v_{3}$ are found to be associated with the initial eccentricities of the participating nucleons, where the values of anisotropic flow coefficients can change when the nuclear distribution is modified [6, 4, 5]. In addition, in Xe–Xe collisions, an observation of a larger value of $v_{2}$ in the most central collisions, as compared to that observed in Pb–Pb collisions, is anticipated due to a deformed nuclear structure of Xe nucleus [7, 8, 10, 9]. In a similar line, the nuclear deformation of Uranium, being a prolate shape nucleus, is reflected in a higher $v_{2}$, when compared to that observed in the most central collisions of Au nuclei, which is close to a spherical nucleus [11]. Similarly, one notices that the values of $v_{2}$ and $v_{3}$ vary with the change in the deformation parameters of Uranium in U–U collisions, and the effect is enhanced for $v_{2}$, as compared to $v_{3}$ [4, 5, 12].
@@
Figure 1: Pictorial representation of an $\alpha$-clustered (left) and non-clustered (right) distribution of nucleons inside the 16O nucleus.
@@
Elliptic flow is one of the observables sensitive to the initial state effects in heavy-ion collisions [13]. However, the value of elliptic flow fluctuates from one event to another, which is a reflection of the fluctuation in the initial energy density and event-by-event participant nucleon distribution before the formation of QGP [14]. These fluctuations in the initial density profile of the participating nucleons can generate odd flow harmonics and lead to fluctuations in the harmonic flow coefficients [14]. Further, the fluctuations in the initial energy density can develop symmetry planes of different orders in various kinematic regions [15, 16, 17]. In other words, the initial density fluctuations can develop transverse momentum ($p_{\rm T}$) and pseudorapidity ($\eta$) dependence of $\psi_{n}$. Therefore, to comprehend the event-by-event initial density fluctuations and their impact on medium evolution, the study of anisotropic flow fluctuations becomes crucial. Anisotropic flow fluctuations are well explored in both experimental and theoretical frontiers of heavy-ion collisions [18, 19, 20, 21, 22, 23, 24]. The measurements by the ATLAS experiment hint at the non-Gaussian behavior of the flow fluctuations in Pb–Pb collisions [25]. Consequently, this observation has led to studies constraining their probability distribution functions [26]. However, the anisotropic flow fluctuations in small collision systems are not yet well understood. Moreover, one may conceive that for a small number of participants, the contribution to fluctuations can increase. Finally, if the evolution of anisotropic flow fluctuations has more than one contributor, then the results can vary significantly from the expectations.
In addition to the collective behavior of the QGP, several other signatures for the presence of a QGP medium have been observed in heavy-ion collisions. For the measurement of such signatures, the baseline measurements are obtained from proton-proton ($pp$) collisions, where the formation of a QCD medium is not anticipated. However, recent observations of strangeness enhancement [27], ridge-like structure [29, 28], radial flow effects [30, 31, 32], etc., in high multiplicity $pp$ collisions have highlighted the significance of studying the small collision systems from different aspects. In addition, recent measurements with ALICE confirm the observation of partonic collectivity in $pp$ and p–Pb collisions, where a characteristic grouping (within $\sim 1\sigma$) and splitting (within $\sim 5\sigma$) of elliptic flow for mesons and baryons at intermediate $p_{\rm T}$ is observed [33]. The LHC and RHIC have plans to inject Oxygen (16O) nuclei to perform p–O and O–O collisions [34, 35], which can bridge the multiplicity gap between $pp$, p–Pb and Pb–Pb collisions and may provide insight into the possible formation of QGP droplets in small systems. As already mentioned, the presence of collective transverse expansion of the medium (also known as collectivity) in ultrarelativistic collisions of nuclear matter is regarded as one of the key signatures of QGP. This leads to the appearance of long-range multi-particle azimuthal correlations in the final state. Through many extensive studies, it has been understood that this collective expansion can be modeled via relativistic hydrodynamics with dissipative effects, which is responsible for these long-range correlations observed in heavy-ion collisions. By comparing the hydrodynamics model predictions with experimentally observed azimuthal anisotropy, it confirms that the initial-state effects such as eccentricity, nuclear deformation, and density fluctuations in the colliding species get embedded in the final-state multi-particle correlations through this collective medium evolution and manifest as finite measurable flow coefficients. Based on these observations from heavy-ion collisions, the existence of long-range multi-particle correlations in high multiplicity events in small collision systems can strongly be associated with the formation of an early deconfined and locally thermalized partonic phase in small collision systems such as O–O collisions.
@@
Furthermore, 16O is a stable nucleus having a double magic number and is theorized to possess clusters of $\alpha$ (4He) particles, where the $\alpha$-particles arrange themselves in the corners of a tetrahedron [36, 37, 38, 39, 40, 41, 42]. Figure 1 shows a pictorial representation of the nucleons inside an 16O nucleus for an $\alpha$-clustered structure (left) and non-clustered distribution of nucleons (right). With the RHIC geometry scan program, 16O collisions would be an interesting study which may unveil the inner geometry of 16O nuclei [12, 43, 44, 45]. In recent years, there have been several studies performed to understand the O–O collisions both at RHIC and LHC energies [51, 52, 50, 46, 47, 48, 6, 49, 53]. A few of them investigate the use of hydrodynamical models [54, 55, 56], and Glauber model calculations [46, 47, 48]. Some intend to study the parton energy loss and jet quenching effects in O–O collisions [57, 58]. Interestingly, there have been a few phenomenological studies that aim to establish the effect of the presence of an $\alpha$-cluster nucleus in the final state particle production [51, 52, 50, 60, 61, 62, 63, 6, 59, 64, 65, 66]. A few of these studies employ O–O collisions, while others employ O–Au, C–Au, O–Pb, and Ne–Pb collisions. Nevertheless, the presence of a clustered nuclear structure of 16O is reflected in the final state azimuthal anisotropy when compared to a Woods-Saxon nuclear density profile. The lattice QCD estimate of a minimum initial energy density requirement of 1 GeV/$\rm fm^{3}$ to form a deconfined partonic medium is conjectured to be fulfilled in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV across all centrality classes as shown in Ref. [49]. In Ref. [59], using a multi-phase transport model (AMPT) [67], authors have shown that an away side broadening is observed in the two-particle correlation function for an $\alpha$-clustered nucleus. In addition, an increased value of $v_{3}$, and $v_{3}/v_{2}$ for most central O–O collisions are also reported for an $\alpha$-clustered structure. A similar observation was reported in Ref. [51], where authors study the anisotropic flow in C–Au collisions with three $\alpha$-clusters of 12C nucleus. However, so far, no studies have shown the effect of change in nuclear density profile on event-by-event flow fluctuations. In this paper, for the first time, we report the behavior of azimuthal anisotropy and flow fluctuations in O–O collisions at $\sqrt{s_{\rm NN}}=7$ TeV using a hydrodynamically expanding system. For this study, we use a hybrid of IPGlasma initial condition, MUSIC hydrodynamics, iSS as a particlisation, and UrQMD as an afterburner for the hadronic cascade. Here, we compare results from an $\alpha$-clustered nuclear distribution with a Woods-Saxon nuclear distribution.
@@ II.1 Simulation framework
In this work, the simulation framework implements the impact-parameter-dependent Glasma (IP-Glasma) initial conditions, MUSIC relativistic viscous hydrodynamics, particle sampling using the iSS package, and performs hadronic cascade through the UrQMD transport model. This hybrid model provides a good description of the particle production and flow in heavy-ion collisions [68]. The key aspects of these models are also discussed briefly below, including the parameters and settings necessary to reproduce the results reported in this study.
@@ II.1.1 IP-Glasma: Initial condition
The IP-Glasma model is used to describe the dynamics of the gluon fields during the collisions of nuclear matter at relativistic speeds [69, 70]. It provides the initial energy-momentum tensor of the classical Yang-Mills (CYM) color Glasma fields ($T_{\rm YM}^{\mu\nu}$), which can be further evolved using relativistic hydrodynamics. This model is based on the color glass condensate effective field theory [71, 72] and uses a classical description of gluon production. This model includes fluctuations in the distributions of nucleons in the nuclear wave functions and also fluctuations in the color charge distributions inside the nucleons. The color charge fluctuations are modeled via the impact-parameter-dependent dipole saturation model (IP-Sat) [73, 74]. For the sub-nucleonic fluctuations, three hot spots are introduced in the proton thickness function, and each hot spot is parametrized using a Gaussian distribution in the transverse plane. The spatial positions of the nucleons inside the colliding nuclei are obtained from the Woods-Saxon distribution function as given below,
@@
Here, $\rho(r)$ is the nuclear charge density at a radial distance $r$ from the center of the nucleus. $r_{0}$, $w$, and $a$ are the mean radius of the nucleus, the nuclear deformation parameter, and nuclear skin depth, respectively. For 16O nucleus, $r_{0}=2.608$ fm, $a=0.513$ fm, $w=-0.051$ [75]. In addition, the exotic $\alpha$–cluster geometry for the 16O nucleus has also been implemented in this study. Four $\alpha$–particles ($2p,~{}2n$) are placed at the vertices of a regular tetrahedron with side length 3.42 fm, which makes the root mean square (RMS) radius of 16O nucleus 2.699 fm [6, 49, 50]. The spatial positions of the nucleons inside an $\alpha$–particle are sampled using the Woods-Saxon distribution given in Eq. (2), with the parameters chosen for the 4He nucleus, which are $r_{0}=0.964$ fm, $w=0.517$ and $a=0.322$ fm. This gives the RMS radius of 1.676 fm for the $\alpha$–particle. A minimum separation distance of 0.4 fm between the nucleons is imposed. Finally, to randomize each nucleus, the entire tetrahedral configuration is rotated through its center randomly in the polar and azimuthal coordinates before its initialization in IP-Glasma. Figure 2 shows the nucleon distribution of 16O nuclei in the x-y plane following $\alpha$-cluster (left) and Woods-Saxon (right) nuclear density profiles. In Fig. 2, for the $\alpha$-cluster nuclear density profile of 16O nuclei, the tetrahedral configuration is not rotated for better visualization of the clustered geometry.
Figure 2: The nucleon distribution of 16O in the $xy$ plane following $\alpha$-cluster (upper) and Woods-Saxon (lower) nuclear density profiles plotted with one million nuclei in each case.
@@ II.1.2 MUSIC: Hydrodynamics
Under the assumption of system under local thermal equilibrium, the initial energy-momentum tensor obtained from the IP-Glasma model at thermalization time $\tau_{0}=0.4$ fm is evolved using the MUSIC framework of relativistic hydrodynamics following the conservation law $\partial_{\mu}T^{\mu\nu}=0$ [78, 79, 80]. The energy-momentum tensor is composed of both ideal and dissipative parts, which are combined to give $T^{\mu\nu}=eu^{\mu}u^{\nu}-(P+\Pi)\Delta^{\mu\nu}+\pi^{\mu\nu}$. Here, $e$ is the energy density, $u^{\mu}$ is the fluid flow four-velocity, $P$ is the thermodynamic pressure, $\Pi$ is the bulk viscous pressure, and $\pi^{\mu\nu}$ is the shear viscous tensor. The projection operator is defined as $\Delta^{\mu\nu}=g^{\mu\nu}-u^{\mu}u^{\nu}$. In this study, we use the boost invariant picture of hydrodynamics and, hence, perform a (2+1)-dimensional evolution. The equation-of-state (EoS) parametrization "s95p-v1.2" is used to close the system of equations in hydrodynamics. This EoS is obtained from the interface of lattice data and hadron resonance gas [81]. We use a fixed $\eta/s=0.12$, and a temperature dependent $\zeta/s(T)$ as described in Ref. [56]. In MUSIC, the hydrodynamic equations are solved using the Kurganov-Tadmor (KT) algorithm [82, 83].